# Model of incompressible turbulent flows via a kinetic theory

### v1

@@ Abstract
Kinetic theory offers a promising alternative to conventional turbulence modelling by providing a mesoscopic perspective that naturally captures non-equilibrium physics such as non-Newtonian effects. In this work, we present an extension and theoretical analysis of the recent kinetic model for incompressible turbulent flows developed by Chen et al. (Atmos. 14(7), 1109, 2023), constructed for unbounded flows. The first extension is to reselect a relaxation time such that the turbulent transport coefficients are obtained more consistently and better align with well-established turbulence theory. The Chapman-Enskog (CE) analysis of the kinetic model reproduces the traditional linear eddy viscosity and gradient diffusion models for Reynolds stress and turbulent kinetic energy flux at the first order, and yields nonlinear eddy viscosity and closure models at the second order. Particularly, a previously unreported CE solution for turbulent kinetic energy flux is obtained. The second extension is to enable the model for wall-bounded turbulent flows with preserved near-wall asymptotic behaviours. This involves developing a low-Reynolds-number kinetic model incorporating wall damping effects and viscous diffusion, with boundary conditions enabling both viscous sublayer resolution and wall function application. Comprehensive validation against experimental and DNS data for turbulent plane Couette flow demonstrates excellent agreement in predicting mean velocity profiles, skin friction coefficients, and Reynolds stress distributions. It reveals that an averaged turbulent flow behaves similarly to a rarefied gas flow at a finite Knudsen number, capturing non-Newtonian effects inaccessible to linear eddy viscosity models. This kinetic model provides a physics-based foundation for turbulence modelling with reduced empirical dependence.
@@ 1 Introduction
Another strategy is to develop kinetic models for turbulence directly from the Boltzmann equation by systematically eliminating small-scale fluctuations through successive averaging procedures. chen1999analysis applied renormalization group techniques to the Boltzmann-BGK equation (bhatnagar1954model) to develop a subgrid-scale viscosity model (SGS). Subsequently, ansumali2004kinetic constructed a turbulence model using a mean-field approach to filter SGS effects, establishing connections with the Smagorinsky-type closures. girimaji2007boltzmann further formulated a Boltzmann kinetic equation by incorporating SGS closure to characterize filtered turbulent flows, applicable to both continuum and non-continuum effects. Recently, yang2025wave developed the wave-particle turbulence simulation method, which conjectured and numerically showed that kinetic representations may be advantageous in describing turbulent non-equilibrium physics. However, a fundamental limitation arises because the averaged equilibrium distribution retains the thermodynamic temperature in the same functional role as turbulent kinetic energy. This prevents the systematic removal of molecular thermal effects through successive averaging processes, resulting in eddy viscosity formulations that exhibit unphysical dependence on molecular temperature rather than being determined solely by turbulent flow characteristics. This limitation highlights fundamental difficulties in applying the conventional Boltzmann equation to turbulence modelling and suggests the need for alternative approaches to statistical averaging.
@@
lundgren1967distribution derived an infinite hierarchy of equations for the multi-point velocity distribution function (VDF) of fluid elements, starting from the incompressible NS equations. In this hierarchy, the $n$th equation contains an unknown ($n+1$)-point VDF, closely resembling the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) (cohen1962fundamental) hierarchy for multi-particle VDF in kinetic theory. The Boltzmann equation for short-range interacting particles and the Fokker-Planck equation for long-range Coulomb interactions are derived from the BBGKY hierarchy through appropriate closure approximations. In subsequent work, lundgren1969model attempted to close the system at the one-point level by employing a relaxation model identical in form to the BGK model. The resulting equation was solved for several idealized problems without solid boundaries, and good agreement with available experiments was found. srinivasan1977turbulent extended the work of lundgren1969model to turbulent flows bounded by solid walls, developing two types of boundary conditions based on the wall function method and employing the discrete ordinate method for numerical solution. However, due to the restriction to regions outside the viscous sublayer, the near-wall Reynolds shear stress was overpredicted, leading to an incorrect shear-stress profile. Lundgren’s model equation provides a very good description of turbulence for the flow situations considered and offers an analytical tool for further study of more complex turbulent flows; yet to our knowledge, this approach has not been pursued in subsequent turbulence research. chen2023average; chen2024average take a different approach but arrives at a similar BGK-type model for the VDF. Their analysis is developed from the idea of Klimontovich-type kinetic equation (klimontovich1969statistical) for fluid elements, which exactly satisfies the incompressible NS equations. Averaging the Klimontovich equation leads to a BGK-type equation without suffering from the mixing of dynamic scales with the thermodynamic temperature. Unlike the model of lundgren1969model, the dissipation rate is incorporated into the collision term rather than added through a source term, providing a more physically realistic representation of dissipation in eddy collisions. These studies provide a rigorous kinetic approach for turbulence modelling, moving beyond heuristic analogies to offer fundamental insights into turbulent fluctuation dynamics and their macroscopic effects. Although the kinetic model developed by chen2023average provides a rigorous theoretical foundation for turbulence modelling, two critical limitations remain. First, the CE analysis for the Reynolds stress is based on the results of the BGK model that does not properly incorporate dissipation of turbulent kinetic energy (chen2004expanded), causing the relaxation time estimation by chen2024average to produce unrealistically high turbulent Prandtl number and higher-order transport coefficients. Second, the model was constructed for unbounded fully developed turbulent flows, and suitable near-wall treatment for wall-bounded turbulence remains to be developed.
@@ 2.1 Kinetic Representation of incompressible turbulence
and $\int fd\bm{\xi}=1$ for an incompressible flow. Taking the zeroth and first order moments of (1), we obtain
@@
which exactly reproduces the NS equations at the hydrodynamic level.
@@
where $F\equiv F(\bm{x},\bm{\xi},t)=\langle f(\bm{x},\bm{\xi},t\rangle$ is an ensemble averaged VDF, with $\langle\cdot\rangle$ denoting the ensemble average. $\mathcal{C}\equiv-\nabla_{\bm{\xi}}\cdot\langle\bm{a}^{\prime}f^{\prime}\rangle$ is the collision operator, with $\bm{a}^{\prime}$ and $f^{\prime}$ denoting the fluctuating components of $\bm{a}$ and $f$, respectively. According to the linearity and incompressibility, the external force $\bar{\bm{a}}$ is given as
@@
But the collision is not conservative with respect to energy, entailing energy dissipation,
$\frac{1}{2}\int\left(\bm{\xi}-\bm{U}\right)^{2}\mathcal{C}d\bm{\xi}=-\epsilon,$ (7)
where $\epsilon$ is the dissipation rate. Therefore, the ensemble averaged Klimontovich-type equation (4) can be regarded as a Boltzmann equation from the kinetic perspective, describing the transport and collision of eddy.
@@
where $\int Fd\bm{\xi}=1$, corresponding to the normalization of the VDF for incompressible flow.
@@
$\partial_{t}K+\left(\bm{U}\cdot\nabla\right)K$ $=-\nabla\cdot\bm{Q}+\bm{\sigma}:\bm{S}-\epsilon,$
where $\bm{S}=\frac{1}{2}\left[\nabla\bm{U}+(\nabla\bm{U})^{T}\right]$ is the rate of strain tensor. It can be seen that $K$, $\bm{\sigma}$, and $\bm{Q}$ are completely determined by the solution of the averaged kinetic equation without any assumptions in this kinetic representation.
@@ 2.2 BGK model for average turbulence dynamics
Since the collision term lacks an explicit form in terms of the averaged variables, one needs to find an appropriate closure in the above kinetic representation. Interestingly, the asymptotic limit of the collision process $\mathcal{C}$ assumes a Gaussian form based on semi-theoretical arguments (chen2023average), which is also consistent with experimental observations in homogeneous turbulent flows where the one-point VDF exhibits a local Gaussian form (tavoularis1981experiments; sreenivasan2021dynamics; sreenivasan2024saturation). As a result, chen2023average modeled the collision term $\mathcal{C}$ as a relaxation process toward the local Gaussian equilibrium with a relaxation time $\tau$, namely, a BGK form,
$\mathcal{C}(F)=\frac{F^{eq}-F}{\tau},$ (10)
where $F^{eq}$ is the equilibrium VDF,
@@
where $K^{eq}$ is the equilibrium TKE. Compared with the canonical Maxwell equilibrium distribution, TKE assumes a role analogous to temperature in classical molecular thermodynamics. In fact, $F^{eq}$ represents a Gaussian equilibrium distribution that excludes the non-trivial flow-induced fluctuations present in $F$; the latter possesses non-Gaussian properties (frisch1996turbulence) such as flatness and skewness.
@@
$\int F^{eq}d\bm{\xi}$ $=\int Fd\bm{\xi}=1,$ (12)
$\int\bm{\xi}F^{eq}d\bm{\xi}$ $=\int\bm{\xi}Fd\bm{\xi}=\bm{U},$
and
@@
and we have the equilibrium TKE $K^{eq}=K-\tau\epsilon$ according to (13).
@@
This choice yields the same viscosity as derived by chen2024average, while providing more reasonable turbulent Prandtl number and improved higher-order transport coefficients. This coefficient is also very close to the value $1/6$ suggested by lundgren1969model, who assumed the relaxation time to scale as $\tau\sim L/\sqrt{K}$ where $L$ is the integral length scale of turbulence (defined through $\epsilon=K^{3/2}/L$).
In the end, the dissipation rate $\epsilon$ remains an unspecified quantity, requiring an additional scale equation to determine its evolution. The most straightforward approach is to adopt an established formulation from conventional turbulence models (e.g., the $K-\epsilon$ model) to prescribe the evolution equation for $\epsilon$. Alternatively, the conventional $K-\omega$ turbulence models can also be utilized, where the relaxation time is expressed as $\tau\sim 1/\omega$, with $\omega\sim\epsilon/K$ being the specific dissipation rate (wilcox1998turbulence). However, these formulations remain incomplete from the perspective of kinetic theory, as they still depend on externally specified parameters. Therefore, a fully self-consistent kinetic theoretic description for inhomogeneous statistical turbulence would be an ultimate goal.
@@ 3.1 Chapman-Enskog analysis
As mentioned in the Introduction, almost all traditional turbulence closures, including the eddy viscosity and gradient diffusion models, aim to approximate the Reynolds stress $\bm{\sigma}$ and the TKE flux $\bm{Q}$ by the mean variables $\bm{U}$ and $K$ in the RANS equations. This hydrodynamic description can be derived by looking for a normal solution to the kinetic equation (4) with the BGK collision term (10) by means of the CE expansion technique (chapman1990mathematical). We first use a reference velocity $U_{ref}$, reference collision time $\tau_{ref}$ and reference length $L_{ref}$ to non-dimensionalise the kinetic equation (4),
@@
$K$ $=K^{(0)}+\mathscr{K}K^{(1)}+\mathscr{K}^{2}K^{(2)}+\cdots,$ (19)
$\bm{\sigma}$ $=\bm{\sigma}^{(0)}+\mathscr{K}\bm{\sigma}^{(1)}+\mathscr{K}^{2}\bm{\sigma}^{(2)}+\cdots,$
$\bm{Q}$ $=\bm{Q}^{(0)}+\mathscr{K}\bm{Q}^{(1)}+\mathscr{K}^{2}\bm{Q}^{(2)}+\cdots.$
@@
$K^{(0)}$ $=K^{eq}\bm{,}$ (20)
$\bm{\sigma}^{(0)}$ $=-\frac{2}{3}K^{eq}\bm{I},$
$\bm{Q}^{(0)}$ $=0,$
@@
$\bm{\sigma}^{(1)}$ $=-\frac{2}{3}K^{(1)}\bm{I}+\frac{4\tau K^{eq}}{3}\bm{S}^{(1)},$ (21)
$\bm{Q}^{(1)}$ $=-\frac{10\tau K^{eq}}{9}\nabla_{1}K^{eq}.$
@@
$\bm{\sigma}$ $\approx\bm{\sigma}^{(0)}+\mathscr{K}\bm{\sigma}^{(1)}=-\frac{2K}{3}\bm{I}+2\nu_{T}\bm{S},$ (22)
$\bm{Q}$ $\approx\bm{Q}^{(0)}+\mathscr{K}\bm{Q}^{(1)}=-\frac{\nu_{T}}{\text{Pr}_{T}}\nabla K,$

### v2

@@ Abstract
Kinetic theory offers a promising alternative to conventional turbulence modelling by providing a mesoscopic perspective that naturally captures non-equilibrium physics such as non-Newtonian effects. In this work, we present an extension and theoretical analysis of the kinetic model for incompressible turbulent flows developed by Chen et al. (Atmos. 14(7), 1109, 2023), constructed for unbounded flows. The first extension is to reselect a relaxation time such that the turbulent transport coefficients are obtained consistently and better align with well-established turbulence theory. The Chapman-Enskog (CE) analysis of the kinetic model reproduces the linear eddy viscosity and gradient diffusion models for Reynolds stress and turbulent kinetic energy flux at the first order, and yields nonlinear eddy viscosity and closure models at the second order. Particularly, a previously unreported CE solution for turbulent kinetic energy flux is obtained. The second extension is to enable the model for wall-bounded turbulent flows with preserved near-wall asymptotic behaviours. This involves developing a low-Reynolds-number model incorporating wall damping effects and viscous diffusion, with boundary conditions enabling both viscous sublayer resolution and wall function application. Comprehensive validation against experimental and DNS data for turbulent Couette flow demonstrates excellent agreement in predicting mean velocity profiles, skin friction coefficients, and Reynolds shear stress distributions, although the near-wall normal stress anisotropy is underestimated. The results show that averaged turbulent flow behaves similarly to rarefied gas flow at finite Knudsen number, capturing non-Newtonian effects beyond linear eddy viscosity models. This kinetic model provides a physics-based foundation for turbulence modelling with reduced empirical dependence.
@@ 1 Introduction
Another strategy is to develop kinetic models for turbulence directly from the Boltzmann equation by systematically eliminating small-scale fluctuations through successive averaging procedures. chen1999analysis applied renormalization group techniques to the Boltzmann-BGK equation (bhatnagar1954model) to develop a subgrid-scale viscosity model (SGS). Subsequently, ansumali2004kinetic constructed a turbulence model using a mean-field approach to filter SGS effects, establishing connections with the Smagorinsky-type closures. However, a fundamental limitation arises because the averaged equilibrium distribution retains the thermodynamic temperature in the same functional role as turbulent kinetic energy. This prevents the systematic removal of molecular thermal effects through successive averaging processes, resulting in eddy viscosity formulations that exhibit unphysical dependence on molecular temperature rather than being determined solely by turbulent flow characteristics. girimaji2007boltzmann rigorously derives a filtered-turbulence kinetic equation from the Boltzmann–BGK framework, in which turbulence effects (Reynolds stress) enter explicitly via additional drift terms. Nevertheless, the resulting Reynolds stress contributions cannot be obtained directly from the distribution function and therefore remain unclosed; in their LES implementation, a conventional Smagorinsky-type SGS closure is employed. This limitation highlights fundamental difficulties in applying the conventional Boltzmann equation to turbulence modelling and suggests the need for alternative approaches to statistical averaging. Recently, yang2025wave developed the wave-particle turbulence simulation method, which conjectured and numerically showed that kinetic representations may be advantageous in describing turbulent non-equilibrium physics.
@@
Several kinetic-based approaches reformulate incompressible NS turbulence in terms of velocity distribution functions (VDFs), providing a complementary viewpoint to conventional RANS. lundgren1967distribution derived an infinite hierarchy of equations for the multi-point VDF of fluid elements, starting from the incompressible NS equations. In this hierarchy, the $n$th equation contains an unknown ($n+1$)-point VDF, closely resembling the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) (cohen1962fundamental) hierarchy for multi-particle VDF in kinetic theory. In subsequent work, lundgren1969model attempted to close the system at the one-point level by employing a relaxation model identical in form to the BGK model. The resulting equation was solved for several idealized problems without solid boundaries, and good agreement with available experiments was found. srinivasan1977turbulent extended the work of lundgren1969model to turbulent flows bounded by solid walls, developing two types of boundary conditions based on the wall function method and employing the discrete ordinate method for numerical solution. However, due to the restriction to regions outside the viscous sublayer, the near-wall Reynolds shear stress was overpredicted, leading to an incorrect shear-stress profile. Lundgren’s model equation provides a very good description of turbulence for the flow situations considered and offers an analytical tool for further study of more complex turbulent flows; yet to our knowledge, this approach has not been pursued in subsequent turbulence research. chen2023average; chen2024average take a different approach but arrive at a similar BGK-type model for the VDF. Their analysis is developed from the idea of Klimontovich-type kinetic equation (klimontovich1969statistical) for fluid elements, which exactly satisfies the incompressible NS equations. Ensemble-averaging the Klimontovich-type equation yields a one-point kinetic equation consistent with lundgren1967distribution and pope1983lagrangian, in which the collision term remains unclosed and needs to be modelled. chen2023average adopt a BGK-type approximation as a closure for this term, while avoiding conflating turbulent fluctuations with the thermodynamic temperature. Unlike the model of lundgren1969model, the dissipation rate is incorporated into the collision term rather than added through a source term, providing a more physically realistic representation of dissipation in eddy collisions. Overall, kinetic-based formulations provide a systematic kinetic framework for turbulence modelling, offering an alternative viewpoint to purely heuristic analogies and facilitating a connection between fluctuation statistics and macroscopic effects. Although the kinetic model developed by chen2023average provides a rigorous theoretical foundation for turbulence modelling, two critical limitations remain. First, the CE analysis for the Reynolds stress is based on the results of the BGK model that does not properly incorporate dissipation rate(chen2004expanded), causing the relaxation time estimation by chen2024average to produce unrealistically high turbulent Prandtl number and higher-order transport coefficients. Second, the model was constructed for unbounded fully developed turbulent flows, and suitable near-wall treatments for wall-bounded turbulence remains to be developed.
@@ 2.1 Kinetic Representation of incompressible turbulence
and $\int fd\bm{\xi}=1$ for an incompressible flow. Taking the moments of (1), we obtain
@@
which exactly reproduces the NS equations at the hydrodynamic level. The Klimontovich-type equation describes representative fluid elements that follow exactly the NS fluid streamlines. Physically, it amounts to a Lagrangian representation of the Navier-Stokes flow. It is in this sense that it is exact and from the first principles.
@@
where $F\equiv F(\bm{x},\bm{\xi},t)=\langle f(\bm{x},\bm{\xi},t\rangle$ is an ensemble averaged VDF, with $\langle\cdot\rangle$ denoting the ensemble average. $\mathcal{C}\equiv-\nabla_{\bm{\xi}}\cdot\langle\bm{a}^{\prime}f^{\prime}\rangle$ is the collision operator, with $\bm{a}^{\prime}$ and $f^{\prime}$ denoting the fluctuating components of $\bm{a}$ and $f$, respectively. And the external force $\bar{\bm{a}}$ is given as
@@
But the collision is not conservative with respect to energy,
$\frac{1}{2}\int\left(\bm{\xi}-\bm{U}\right)^{2}\mathcal{C}d\bm{\xi}=-\nabla\cdot\langle p^{\prime}\bm{u}^{\prime}\rangle+\nu_{0}\nabla^{2}K-\epsilon,$ (7)
where $p^{\prime}$ and $\bm{u}^{\prime}$ denoting the fluctuating components of $p$ and $\bm{u}$, respectively. $\epsilon=\nu_{0}\langle\nabla\bm{u}^{\prime}:\nabla\bm{u}^{\prime}\rangle$ is the dissipation rate. The kinetic equation obtained by ensemble averaging the Klimontovich-type equation is also exact before any closure model is applied for the collision term.
@@
where $\int Fd\bm{\xi}=1$, corresponding to the normalization of the VDF for incompressible flow. More generally, arbitrary higher-order central moments of the velocity fluctuations can be obtained from the VDF as $\langle u^{\prime}_{i_{1}}u^{\prime}_{i_{2}}\dots u^{\prime}_{i_{n}}\rangle=\int(\xi_{i_{1}}-U_{i_{1}})(\xi_{i_{2}}-U_{i_{2}})\dots(\xi_{i_{n}}-U_{i_{n}})Fd\bm{\xi}$.
@@
$\partial_{t}K+\left(\bm{U}\cdot\nabla\right)K$ $=-\nabla\cdot\bm{Q}+\bm{\sigma}:\bm{S}-\nabla\cdot\langle p^{\prime}\bm{u}^{\prime}\rangle+\nu_{0}\nabla^{2}K-\epsilon,$
where $\bm{S}=\frac{1}{2}\left[\nabla\bm{U}+(\nabla\bm{U})^{T}\right]$ is the rate of strain tensor. It can be seen that $K$, $\bm{\sigma}$, and $\bm{Q}$ are completely determined by the solution of the averaged kinetic equation without any assumptions in this kinetic representation. Moreover, in principle, transport equations for any other higher-order central moments can be derived by taking corresponding moments of the kinetic equation.
@@ 2.2 BGK model for average turbulence dynamics
Since the collision term lacks an explicit form in terms of the averaged variables, one needs to find an appropriate closure in the above kinetic representation. Interestingly, the asymptotic limit of the collision process $\mathcal{C}$ assumes a Gaussian form based on semi-theoretical arguments (chen2023average), which is also consistent with experimental observations in homogeneous turbulent flows where the one-point VDF exhibits a local Gaussian form (tavoularis1981experiments; sreenivasan2021dynamics; sreenivasan2024saturation). As a result, chen2023average modeled the collision term $\mathcal{C}$ as a relaxation process toward the local Gaussian form with a relaxation time $\tau$, namely, a BGK form,
$\mathcal{C}_{\text{BGK}}(F)=\frac{F^{eq}-F}{\tau},$ (10)
where $F^{eq}$ is the local equilibrium VDF,
@@
where $K^{eq}$ is the local equilibrium TKE. Notably, in dissipative turbulence $F^{eq}$ denotes a target function representing the asymptotic state of the BGK collision process, which is in general never achieved but serves as the direction towards which the collision process drives $F$. It excludes the non-trivial flow-induced fluctuations contained in $F$; the latter exhibits non-Gaussian properties (frisch1996turbulence) such as flatness and skewness. In the homogeneous case, the kinetic equation with the BGK model (10) admits a Gaussian solution whose variance equals the instantaneous TKE.
@@
$\int F^{eq}d\bm{\xi}=\int Fd\bm{\xi}=1,\quad\quad\int\bm{\xi}F^{eq}d\bm{\xi}=\int\bm{\xi}Fd\bm{\xi}=\bm{U}.$ (12)
To simplify the analysis, chen2023average considered statistically homogeneous, high-Reynolds-number turbulence (or equivalently, the spatially averaged TKE budget), for which the divergence-type transport terms in Eq. (7) do not contribute to the net balance,
@@
and we have the equilibrium TKE $K^{eq}=K-\tau\epsilon$ according to (13). In conventional RANS modeling, the pressure-diffusion term $\nabla\cdot(p^{\prime}\bm{u}^{\prime})$ is rarely treated separately; instead it is commonly modeled together with the TKE flux $\bm{Q}$ using a gradient diffusion closure (launder1983numerical; wilcox1998turbulence). In the present work, we neglect this term because it is quite small for simple flows (mansour1988reynolds; pope2001turbulent). Additionally, for low-Reynolds-number turbulence and/or near-wall and inhomogeneous flows, the molecular diffusion $\nu_{0}\nabla^{2}K$ may be non-negligible and will be explicitly retained and discussed in Sec. 4.1.
@@
This choice yields the same viscosity as derived by chen2024average, while providing more reasonable turbulent Prandtl number and improved higher-order transport coefficients.
In the end, the dissipation rate $\epsilon$ remains an unspecified quantity, requiring an additional scale equation to determine its evolution. The most straightforward approach is to adopt an established formulation from conventional turbulence models (e.g., the $K-\epsilon$ model) to prescribe the evolution equation for $\epsilon$. Alternatively, the conventional $K-\omega$ turbulence models can also be utilized, where the relaxation time is expressed as $\tau\sim 1/\omega$, with $\omega\sim\epsilon/K$ being the specific dissipation rate (wilcox1998turbulence). These formulations remain incomplete from the perspective of kinetic theory, as they still depend on externally specified parameters. Therefore, a fully self-consistent kinetic theoretic description for inhomogeneous statistical turbulence would be an ultimate goal.
@@ 3.1 Chapman-Enskog analysis
As mentioned in the Introduction, almost all traditional turbulence closures, including the eddy viscosity and gradient diffusion models, aim to approximate the Reynolds stress $\bm{\sigma}$ and the TKE flux $\bm{Q}$ by the mean variables $\bm{U}$ and $K$ in the RANS equations. This hydrodynamic description can be derived by looking for a normal solution to the kinetic equation (4) with the BGK collision term (10) by means of the CE expansion technique (chapman1990mathematical). It should be noted that the CE expansion is primarily used to make the relationship with traditional closure models transparent. We first use a reference velocity $U_{ref}$, reference collision time $\tau_{ref}$ and reference length $L_{ref}$ to non-dimensionalise the kinetic equation (4),
@@
$K=\sum_{n\geq 0}\mathscr{K}^{n}K^{(n)},\qquad\bm{\sigma}=\sum_{n\geq 0}\mathscr{K}^{n}\bm{\sigma}^{(n)},\qquad\bm{Q}=\sum_{n\geq 0}\mathscr{K}^{n}\bm{Q}^{(n)},$ (19)
@@
$K^{(0)}=K^{eq}\bm{,}\quad\bm{\sigma}^{(0)}=-\frac{2}{3}K^{eq}\bm{I},\quad\bm{Q}^{(0)}=0,$ (20)
@@
$\bm{\sigma}^{(1)}=-\frac{2}{3}K^{(1)}\bm{I}+\frac{4\tau K^{eq}}{3}\bm{S}^{(1)},\quad\bm{Q}^{(1)}=-\frac{10\tau K^{eq}}{9}\nabla_{1}K^{eq}.$ (21)
@@
$\bm{\sigma}\approx\bm{\sigma}^{(0)}+\mathscr{K}\bm{\sigma}^{(1)}=-\frac{2K}{3}\bm{I}+2\nu_{T}\bm{S},\quad\bm{Q}\approx\bm{Q}^{(0)}+\mathscr{K}\bm{Q}^{(1)}=-\frac{\nu_{T}}{\text{Pr}_{T}}\nabla K,$ (22)